# QiMeng-CRUX: Narrowing the Gap between Natural Language and Verilog via Core Refined Understanding eXpression

### v2

@@ Abstract
Large language models (LLMs) have shown promising capabilities in hardware description language (HDL) generation. However, existing approaches often rely on free-form natural language descriptions that are often ambiguous, redundant, and unstructured, which poses significant challenges for downstream Verilog code generation. We treat hardware code generation as a complex transformation from an open-ended natural language space to a domain-specific, highly constrained target space. To bridge this gap, we introduce Core Refined Understanding eXpression (CRUX), a structured intermediate space that captures the essential semantics of user intent while organizing the expression for precise Verilog code generation. We further design a two-stage training framework, comprising Joint Expression Modeling and Dual-Space Optimization, to enhance the quality of both CRUX and Verilog code. Experiments across multiple Verilog generation benchmarks demonstrate that our model, QiMeng-CRUX-V, achieves state-of-the-art performance among general models, particularly under challenging design tasks. Furthermore, the CRUX space proves transferable and beneficial when used as input prompts for other code models, highlighting its effectiveness in narrowing the gap between free-form natural language descriptions and precise Verilog generation.
@@
Figure 2: An example from the VerilogEval-v2 Spec-to-RTL (verilogeval-v2) benchmark shows that the original description lacks explicit emphasis on critical design details. Feeding this directly to LLMs often leads to incorrect implementations due to misinterpretation of key details. While the model possesses the capability to generate correct code, the ambiguity and underspecification in the input hinder accurate realization. In contrast, providing the model with CRUX enables more reliable alignment with the desired design specification.
@@ Introduction
Large language models (LLMs) have shown remarkable progress in code generation for general-purpose programming languages, spurring increasing interest in automating hardware description language (HDL) code generation, such as Verilog (dawn). Recent efforts have substantially improved LLM performance on hardware design tasks through advancements in many aspects, such as data construction (rtlcoder; codev; mg-verilog), feedback mechanisms (insightsfromverification; origen; codev-r1), and training paradigms (aotovcoder; betterv). However, most existing LLM-based methods for Verilog generation focus primarily on the code generation phase and rely solely on free-form natural language descriptions, with limited attention to how the expression of descriptions influences downstream performance.
In practice, free-form natural language descriptions of hardware design from users often suffer from ambiguity in intent, redundancy in expression, and irregularity in structure, particularly in critical modules involving sequential logic, finite state machines (FSMs), and complex control behaviors (craftrtl). Thus, crucial information is often scattered, underspecified, or only implicitly conveyed through the text, posing significant challenges for the model to organize and understand the necessary semantics. In contrast, Verilog is a formal specification language for modeling hardware circuits, characterized by strict structural constraints, strong semantic rigor, and high demands for domain expertise (betterv). Such gap reveals a fundamental challenge in Verilog code generation task: transforming from a highly expressive and inherently unstructured description space into a structured code space governed by syntactic rigidity, modular hierarchy, and functional alignment. The high difficulty of this transformation poses significant obstacles to existing LLM-based approaches, often resulting in misaligned design intent, semantic drift, and distorted implementations in engineering.
Motivated by this gap, we investigate the effect of input formulation and find semantically structured and refined expressions leading to better generation performance. As illustrated in Figure 1, even when the descriptions of the two tasks contain the same underlying information, the model’s performance can vary substantially depending on how the information is expressed. Notably, when the original input of Spec-to-RTL (verilogeval-v2) is restructured to expose its key components more explicitly and precisely, the models demonstrate significant improvements in generation accuracy, suggesting the importance of input clarity and organization in guiding LLMs toward robust code generation.
@@
Comprehensive experiments on multiple Verilog generation benchmarks demonstrate that our model QiMeng-CRUX-V, achieves state-of-the-art performance across both general-purpose and Verilog-specific models, especially under challenging and realistic design scenarios like Spec-to-RTL(from 49.3% to 64.7% with T=0, from 46.8% to 64.4% with T=0.8) and RTLLM-v2(from 50.9% to 63.8%). In addition, the CRUX space proves to be a robust and semantically meaningful intermediate guidance that enhances performance even when used as a prompt for other code models without any further training. These findings highlight the potential of CRUX as a powerful and effective approach for supporting precise Verilog code generation. Code is available at https://github.com/Taskii-Lei/QiMeng-CRUX-V.
@@
High-quality datasets play a central role in Verilog code generation. Early works such as Dave (pearce2020dave) explored fine-tuning of GPT-2 to generate Verilog snippets from natural language, marking one of the first attempts to align LLMs with hardware tasks. Recently, several works mined codes from public sources and refined task descriptions (rtlcoder; verigen; codev). CraftRTL (craftrtl) extended this by introducing non-text tasks like FSM to increase diversity. Other efforts including hdl2v (hdl2v) and MG-Verilog (mg-verilog), focused on improving data quality through code translation and fine-grained annotation. These datasets collectively enhance both the scale and quality of Verilog training data.
@@
Given that Verilog code is ultimately executed and verified by downstream EDA tools, recent studies have explored two main strategies for incorporating feedback to improve generation quality. One line of work applies feedback at the inference stage, where either verification feedback (insightsfromverification; automaticallyimproving; romanwasnotbuilt), or model feedback (origen; vrank) is used to identify and revise incorrect outputs. The other line leverages RL-based post-training (schulman2017ppo; bai2022RLHF; shao2024grpo; yu2025dapo) to optimize generation with reward functions (coderl; dou2024stepcoder; veriseek; codev-r1), which may incorporate verification feedback, structural constraints, and design priors. This enables models to move beyond static supervision and align more closely with hardware-centric design objectives.
@@
Motivated by the success of DeepSeek-R1 (guo2025deepseek), which demonstrates the benefits of chain-of-thought (CoT) reasoning with RL, recent studies have tried to incorporate such diagram into Verilog code generation with verification feedback. Most CoT supervision is obtained through distillation from stronger models (verithoughts; reasoningv; codev-r1), while verireason adopts optimization-based pipelines to construct thinking trajectories. While effective, CoT-based methods often require higher training costs and longer inference time compared to general models (haven).
@@
To equip the model with the ability to construct CRUX and generate Verilog code, we begin with supervised fine-tuning (SFT). However, most existing datasets are synthesized and exhibit limited diversity in descriptions. So we augment the dataset of CodeV(codev) by simulating real-world user phrasing via introducing variations, ambiguities, and structural incompleteness. Then we derive corresponding CRUX space to explicitly extract and organize the core design intent, resulting in a reconstructed dataset of triplets (R, X, V), where R denotes the realistic task specifications (RealSpec), X is the CRUX, and V is the target Verilog code implementations. Finally, we perform SFT using both CRUX and Verilog code as learning targets. The model is trained to generate the CRUX first from the RealSpec, and then the Verilog code.
@@
Given the diverse nature of Verilog tasks, we aim to identify which problems intrinsically require structural guidance beyond plain-text prompts. In particular, we are interested in a subset of circuit-specific tasks that cannot be adequately described or resolved through natural language alone, such as Karnaugh map interpretation, FSM modeling, or waveform analysis(craftrtl). To systematically explore this heterogeneity, we categorize the dataset into three types: Easy Questions, Special Non-text and Normal Data. We begin by prompting LLMs with task descriptions to generate corresponding Verilog code. Functional correctness is then automatically verified via testbenches comparing generated outputs with references under identical inputs. Tasks passing verification are labeled Easy Questions. Among failures, those containing keywords such as ”k-map,” ”FSM,” or ”sequential/waveform” are classified as Special Non-Text, with the rest assigned to Normal Data.
@@
We define the CRUX space as comprising three components: Module Interface, Core Functions, and Key Considerations. The Module Interface comes from the reference code header, while the other two are derived differently depending on the corpus type, as illustrated in the left of Figure 4. (i) Easy Questions: The descriptions are considered sufficiently clear and complete, so we directly adopt them as Core Functions, leaving Key Considerations empty. (ii) Normal Data: As the primary source for constructing CRUX, these tasks are processed by prompting DeepSeek-R1 (guo2025deepseek) with the descriptions and reference codes to extract and refine the Core Functions and Key Considerations. (iii) Special Non-Text: We use Qwen2.5-Coder-32B-Instruct (hui2024qwen2) as a circuit parser to generate diagrams and concise analyses from both the description and code. This includes “State → Condition → Next State” for FSMs, logical expressions for K-maps or truth tables, and core behavior for sequential circuits. The generated diagram and code are then validated by Qwen2.5-Coder-32B-Instruct. Only valid samples are retained to construct the RealSpec and CRUX, while parsing failures or incorrect cases are reclassified into Normal Data to process.
@@
After the steps above, we reconstruct the original dataset of CodeV(codev) into (R, X, V) tuples and perform SFT to establish the LLM’s initial capacity for generating CRUX and correct Verilog code based on RealSpec.
@@
Foundation
General Models
GPT-4o 67.7 75.5 77.2 60.1 71.4 74.5 41.7 65.9 -
@@
Foundation
Reasoning Models
Deepseek-R1-671B 81.0 87.4 89.5 81.5 87.6 88.5 64.8 82.9 -
@@
Verilog-Specific
Reasoning Models
HaVen-7B 77.3 81.2 - 61.1 64.8 - 62.2 - -
@@
General
Code Models
Qwen2.5-Coder-7B 50.1 66.5 70.9 22.9 36 39.5 32.2 48.2 56.0
@@
Verilog-Specific
General Models
RTLCoder-6.7B 61.2 76.5 81.8 41.6 50.1 53.4 35.8 40.3 43.1
@@
QiMeng-CRUX-V-SFT-7B(Ours) 76.8 87.3 91.8 63.2 73.1 76.1 49.0 64.1 72.0
QiMeng-CRUX-V-Final-7B(Ours) 82.9 88.3 90.2 65.2 72.0 73.8 62.8 69.0 71.5
@@
Foundation
General Models
GPT-4o 59.0 56.1 57.6 62.5 61.4 64.1 56.5 70.3 75.2
@@
Verilog-Specific
Reasoning Models
HaVen-7B - - - - - 54.6 - - -
@@
General
Code Models
Qwen2.5-Coder-7B 30.1 25.6 30.5 23.0 22.0 31.3 36.1 52.4 57.6
@@
Verilog-Specific
General Models
RTLCoder-6.7B 35.9 31.5 33.7 36.8 30.9 31.1 33.6 45.3 49.2
@@
QiMeng-CRUX-V-SFT-7B(Ours) 64.7 58.9 65.5 59.6 57.0 61.1 52.3 68.6 73.4
QiMeng-CRUX-V-Final-7B(Ours) 68.0 66.7 67.6 64.7 64.4 64.2 63.8 70.6 73.9
@@
Our dataset is derived from the CodeV(codev) corpus, which was constructed by hierarchically parsing Verilog code with GPT to obtain (Description, Verilog Code) pairs. They used the Rouge-L metric to measure the similarity between the dataset and the benchmark, and removed samples with Rouge-L >> 0.5 to avoid data contamination. Following Corpus Categorization, we obtain about 40k Easy Questions, 18k Special Non-Text, and 107k Normal Data samples. During the Interface Degradation process, the complete interface is retained with a probability of 0.2, while all other variation elements are applied independently with a probability of 0.5. Only about 24k samples insert interface information in the middle of the description; the remaining samples include it only at the end.
@@
To ensure a comprehensive evaluation, we adopt several representative and widely used Verilog generation benchmarks, including VerilogEval-V1(verilogeval-v1), VerilogEval-V2(verilogeval-v2), RTLLM-V1(rtllm-v1), and RTLLM-V2(rtllm-v2). VerilogEval-V1 consists of two sets of tasks: Machine (143 tasks automatically generated by GPT) and Human (156 hand-crafted tasks), both primarily aimed at evaluating code completion capabilities. Building upon this, VerilogEval-V2 further enhances the quality and diversity by introducing two subtasks: Code-Completion that refined from V1-Human, and Spec-to-RTL which imposes additional requirements on the model’s ability to understand and generate core functional logic based on the specification. The RTLLM benchmark consists of four categories: Arithmetic, Memory, Control, and Miscellaneous, covering a wide range of challenges, including FSMs, combinational and sequential circuits, and long-text specifications. It poses a higher level of difficulty compared to VerilogEval. RTLLM-V2 expands upon RTLLM-V1, increasing the total number of tasks from 29 to 50, with V1 forming a subset of V2. Since some prior works only report results on RTLLM-V1, we also include our performance on V1 for fair comparison.
@@
We adopt Qwen2.5-Coder-7B-Instruct (hui2024qwen2) as the backbone model and follow the proposed two-stage framework, applying supervised fine-tuning (SFT) for 2 epochs with LLaMaFactory (llamafactory) in Stage I, and CRUX-enhanced GRPO training for 1 epoch with EasyR1 (easyr1) in Stage II.
@@
Notably, the original VerilogEval-V2 paper requires reporting only pass@1 under two settings to highlight practical LLM evaluation: a low-temperature setting (T=0.0, top_p=0.01, n=1) and a high-temperature setting (T=0.8, top_p=0.95, n=20) (verilogeval-v2). However, many prior works still adopt a strategy of selecting the best performance among Temperature = 0.2, 0.5, 0.8. To enable fair comparison, we also include pass@1 results under this strategy, denoted as T-cross. For the other benchmarks, we follow common practice and generate $n=20$ independent solutions per task to compute pass@1, pass@5, and pass@10, reporting the best performance across T=0.2, 0.5, and 0.8.
@@
Our main experimental results are presented in Table 1 and Table 2. We report the performance of Foundation Models (deepseek-v3; qwen3technicalreport), General Code Models (hui2024qwen2; deepseek-coder), and Verilog-Specific General Models (rtlcoder; codev; origen; insightsfromverification). Given the remarkable performance of Reasoning Models in recent years, we separately group Deepseek-R1 (guo2025deepseek) and Verilog-Specific Reasoning Models (haven; codev-r1).
@@
The main results demonstrate our QiMeng-CRUX-V model achieves the best pass@1 performance across all benchmarks within both the General-Purpose Code Models and Verilog-Specific General Models, especially under challenging and realistic scenarios. It significantly outperforms GPT-4o on multiple tasks, most notably on RTLLM-v2, where it achieves a pass@1 of 63.8%, surpassing the previous SOTA by 12.8%, and outperforming Deepseek-V3, reaching performance comparable with Deepseek-R1. On VerilogEval-v2, QiMeng-CRUX-V brings substantial improvements over previous SOTA models across nearly all metrics, such as from 58.3% to 68.0% (+9.7%) in pass@1 on the CC task, and from 49.3% to 64.7% (+15.4%) on the SR task. These results suggest that CRUX offers significant advantages in tackling complex code generation tasks and real-world design inputs.
Plus, QiMeng-CRUX-V demonstrates robust performance across diverse task settings and specification styles. Originating from the CodeV Datasets, QiMeng-CRUX-V consistently performs better across tasks with diverse input forms and exhibits stronger robustness to specification variations than CodeV2.5. It achieves balanced performance across CC and SR (68.0% in CC and 64.7% in SR), avoiding the large gap observed in CodeV2.5 (58.3% in CC and 44.8% in SR). Furthermore, QiMeng-CRUX-V achieves substantial improvements across all metrics on the RTLLM benchmark, highlighting the effectiveness and robustness of the CRUX methodology.

### v3

@@ Abstract
Large language models (LLMs) have shown promising capabilities in hardware description language (HDL) generation. However, existing approaches often rely on free-form natural language descriptions that are often ambiguous, redundant, and unstructured, which poses significant challenges for downstream Verilog code generation. We treat hardware code generation as a complex transformation from an open-ended natural language space to a domain-specific, highly constrained target space. To bridge this gap, we introduce Core Refined Understanding eXpression (CRUX), a structured intermediate space that captures the essential semantics of user intent while organizing the expression for precise Verilog code generation. We further design a two-stage training framework, comprising Joint Expression Modeling and Dual-Space Optimization, to enhance the quality of both CRUX and Verilog code. Experiments across multiple Verilog generation benchmarks demonstrate that our model, QiMeng-CRUX, achieves state-of-the-art performance among general models, particularly under challenging design tasks. Furthermore, the CRUX space proves transferable and beneficial when used as input prompts for other code models, highlighting its effectiveness in narrowing the gap between free-form natural language descriptions and precise Verilog generation.
@@
Figure 2: An example from the VerilogEval-v2 Spec-to-RTL (Pinckney et al. 2025) benchmark shows that the original description lacks explicit emphasis on critical design details. Feeding this directly to LLMs often leads to incorrect implementations due to misinterpretation of key details. While the model possesses the capability to generate correct code, the ambiguity and underspecification in the input hinder accurate realization. In contrast, providing the model with CRUX enables more reliable alignment with the desired design specification.
@@ Introduction
Large language models (LLMs) have shown remarkable progress in code generation for general-purpose programming languages, spurring increasing interest in automating hardware description language (HDL) code generation, such as Verilog (Chen et al. 2024). Recent efforts have substantially improved LLM performance on hardware design tasks through advancements in many aspects, such as data construction (Liu et al. 2024b; Zhao et al. 2025a; Zhang et al. 2024), feedback mechanisms (Wang et al. 2025a; Cui et al. 2024; Zhu et al. 2025), and training paradigms (Gao et al. 2024; Pei et al. 2024). However, most existing LLM-based methods for Verilog generation focus primarily on the code generation phase and rely solely on free-form natural language descriptions, with limited attention to how the expression of descriptions influences downstream performance.
In practice, free-form natural language descriptions of hardware design from users often suffer from ambiguity in intent, redundancy in expression, and irregularity in structure, particularly in critical modules involving sequential logic, finite state machines (FSMs), and complex control behaviors (Liu et al. 2025). Thus, crucial information is often scattered, underspecified, or only implicitly conveyed through the text, posing significant challenges for the model to organize and understand the necessary semantics. In contrast, Verilog is a formal specification language for modeling hardware circuits, characterized by strict structural constraints, strong semantic rigor, and high demands for domain expertise (Pei et al. 2024). Such gap reveals a fundamental challenge in Verilog code generation task: transforming from a highly expressive and inherently unstructured description space into a structured code space governed by syntactic rigidity, modular hierarchy, and functional alignment. The high difficulty of this transformation poses significant obstacles to existing LLM-based approaches, often resulting in misaligned design intent, semantic drift, and distorted implementations in engineering.
Motivated by this gap, we investigate the effect of input formulation and find semantically structured and refined expressions leading to better generation performance. As illustrated in Figure 1, even when the descriptions of the two tasks contain the same underlying information, the model’s performance can vary substantially depending on how the information is expressed. Notably, when the original input of Spec-to-RTL (Pinckney et al. 2025) is restructured to expose its key components more explicitly and precisely, the models demonstrate significant improvements in generation accuracy, suggesting the importance of input clarity and organization in guiding LLMs toward robust code generation.
@@
Comprehensive experiments on multiple Verilog generation benchmarks demonstrate that our model QiMeng-CRUX, achieves state-of-the-art performance across both general-purpose and Verilog-specific models, especially under challenging and realistic design scenarios like Spec-to-RTL (from 49.3% to 64.7% with T=0, from 46.8% to 64.4% with T=0.8) and RTLLM-v2 (from 50.9% to 63.8%). In addition, the CRUX space proves to be a robust and semantically meaningful intermediate guidance that enhances performance even when used as a prompt for other code models without any further training. These findings highlight the potential of CRUX as a powerful and effective approach for supporting precise Verilog code generation. Code is available at https://github.com/Taskii-Lei/QiMeng-CRUX-V.
@@
High-quality datasets play a central role in Verilog code generation. Early works such as Dave (Pearce et al. 2020) explored fine-tuning of GPT-2 to generate Verilog snippets from natural language, marking one of the first attempts to align LLMs with hardware tasks. Recently, several works mined codes from public sources and refined task descriptions (Liu et al. 2024b; Thakur et al. 2024; Zhao et al. 2025a). CraftRTL (Liu et al. 2025) extended this by introducing non-text tasks like FSM to increase diversity. Other efforts including hdl2v (Hong et al. 2025) and MG-Verilog (Zhang et al. 2024), focused on improving data quality through code translation and fine-grained annotation. These datasets collectively enhance both the scale and quality of Verilog training data.
@@
Given that Verilog code is ultimately executed and verified by downstream EDA tools, recent studies have explored two main strategies for incorporating feedback to improve generation quality. One line of work applies feedback at the inference stage, where either verification feedback (Wang et al. 2025a; Blocklove et al. 2025; Nakkab et al. 2024), or model feedback (Cui et al. 2024; Zhao et al. 2025b) is used to identify and revise incorrect outputs. The other line leverages RL-based post-training (Schulman et al. 2017; Bai et al. 2022; Shao et al. 2024; Yu et al. 2025) to optimize generation with reward functions (Le et al. 2022; Dou et al. 2024; Wang et al. 2024; Zhu et al. 2025), which may incorporate verification feedback, structural constraints, and design priors. This enables models to move beyond static supervision and align more closely with hardware-centric design objectives.
@@
Motivated by the success of DeepSeek-R1 (Guo et al. 2025), which demonstrates the benefits of chain-of-thought (CoT) reasoning with RL, recent studies have tried to incorporate such diagram into Verilog code generation with verification feedback. Most CoT supervision is obtained through distillation from stronger models (Yubeaton et al. 2025; Qin et al. 2025; Zhu et al. 2025), while Wang et al. adopts optimization-based pipelines to construct thinking trajectories. While effective, CoT-based methods often require higher training costs and longer inference time compared to general models (Yang et al. 2025).
@@
To equip the model with the ability to construct CRUX and generate Verilog code, we begin with supervised fine-tuning (SFT). However, most existing datasets are synthesized and exhibit limited diversity in descriptions. So we augment the dataset of CodeV (Zhao et al. 2025a) by simulating real-world user phrasing via introducing variations, ambiguities, and structural incompleteness. Then we derive corresponding CRUX space to explicitly extract and organize the core design intent, resulting in a reconstructed dataset of triplets (R, X, V), where R denotes the realistic task specifications (RealSpec), X is the CRUX, and V is the target Verilog code implementations. Finally, we perform SFT using both CRUX and Verilog code as learning targets. The model is trained to generate the CRUX first from the RealSpec, and then the Verilog code.
@@
Given the diverse nature of Verilog tasks, we aim to identify which problems intrinsically require structural guidance beyond plain-text prompts. In particular, we are interested in a subset of circuit-specific tasks that cannot be adequately described or resolved through natural language alone, such as Karnaugh map interpretation, FSM modeling, or waveform analysis (Liu et al. 2025). To systematically explore this heterogeneity, we categorize the dataset into three types: Easy Questions, Special Non-text and Normal Data. We begin by prompting LLMs with task descriptions to generate corresponding Verilog code. Functional correctness is then automatically verified via testbenches comparing generated outputs with references under identical inputs. Tasks passing verification are labeled Easy Questions. Among failures, those containing keywords such as ”k-map,” ”FSM,” or ”sequential/waveform” are classified as Special Non-Text, with the rest assigned to Normal Data.
@@
We define the CRUX space as comprising three components: Module Interface, Core Functions, and Key Considerations. The Module Interface comes from the reference code header, while the other two are derived differently depending on the corpus type, as illustrated in the left of Figure 4. (i) Easy Questions: The descriptions are considered sufficiently clear and complete, so we directly adopt them as Core Functions, leaving Key Considerations empty. (ii) Normal Data: As the primary source for constructing CRUX, these tasks are processed by prompting DeepSeek-R1 (Guo et al. 2025) with the descriptions and reference codes to extract and refine the Core Functions and Key Considerations. (iii) Special Non-Text: We use Qwen2.5-Coder-32B-Instruct (Hui et al. 2024) as a circuit parser to generate diagrams and concise analyses from both the description and code. This includes “State → Condition → Next State” for FSMs, logical expressions for K-maps or truth tables, and core behavior for sequential circuits. The generated diagram and code are then validated by Qwen2.5-Coder-32B-Instruct. Only valid samples are retained to construct the RealSpec and CRUX, while parsing failures or incorrect cases are reclassified into Normal Data to process.
@@
After the steps above, we reconstruct the original dataset of CodeV (Zhao et al. 2025a) into (R, X, V) tuples and perform SFT to establish the LLM’s initial capacity for generating CRUX and correct Verilog code based on RealSpec.
@@
Foundation General Models GPT-4o 67.7 75.5 77.2 60.1 71.4 74.5 41.7 65.9 -
@@
Foundation Reasoning Models Deepseek-R1-671B 81.0 87.4 89.5 81.5 87.6 88.5 64.8 82.9 -
@@
Verilog-Specific Reasoning Models HaVen-7B 77.3 81.2 - 61.1 64.8 - 62.2 - -
@@
General Code Models Qwen2.5-Coder-7B 50.1 66.5 70.9 22.9 36 39.5 32.2 48.2 56.0
@@
Verilog-Specific General Models RTLCoder-6.7B 61.2 76.5 81.8 41.6 50.1 53.4 35.8 40.3 43.1
@@
QiMeng-CRUX-SFT-7B (Ours) 76.8 87.3 91.8 63.2 73.1 76.1 49.0 64.1 72.0
QiMeng-CRUX-Final-7B (Ours) 82.9 88.3 90.2 65.2 72.0 73.8 62.8 69.0 71.5
@@
Foundation General Models GPT-4o 59.0 56.1 57.6 62.5 61.4 64.1 56.5 70.3 75.2
@@
Verilog-Specific Reasoning Models HaVen-7B - - - - - 54.6 - - -
@@
General Code Models Qwen2.5-Coder-7B 30.1 25.6 30.5 23.0 22.0 31.3 36.1 52.4 57.6
@@
Verilog-Specific General Models RTLCoder-6.7B 35.9 31.5 33.7 36.8 30.9 31.1 33.6 45.3 49.2
@@
QiMeng-CRUX-SFT-7B (Ours) 64.7 58.9 65.5 59.6 57.0 61.1 52.3 68.6 73.4
QiMeng-CRUX-Final-7B (Ours) 68.0 66.7 67.6 64.7 64.4 64.2 63.8 70.6 73.9
@@
Our dataset is derived from the CodeV (Zhao et al. 2025a) corpus, which was constructed by hierarchically parsing Verilog code with GPT to obtain (Description, Verilog Code) pairs. They used the Rouge-L metric to measure the similarity between the dataset and the benchmark, and removed samples with Rouge-L >> 0.5 to avoid data contamination. Following Corpus Categorization, we obtain about 40k Easy Questions, 18k Special Non-Text, and 107k Normal Data samples. During the Interface Degradation process, the complete interface is retained with a probability of 0.2, while all other variation elements are applied independently with a probability of 0.5. Only about 24k samples insert interface information in the middle of the description; the remaining samples include it only at the end.
@@
To ensure a comprehensive evaluation, we adopt several representative and widely used Verilog generation benchmarks, including VerilogEval-V1 (Liu et al. 2023), VerilogEval-V2 (Pinckney et al. 2025), RTLLM-V1 (Lu et al. 2024), and RTLLM-V2 (Liu et al. 2024c). VerilogEval-V1 consists of two sets of tasks: Machine (143 tasks automatically generated by GPT) and Human (156 hand-crafted tasks), both primarily aimed at evaluating code completion capabilities. Building upon this, VerilogEval-V2 further enhances the quality and diversity by introducing two subtasks: Code-Completion that refined from V1-Human, and Spec-to-RTL which imposes additional requirements on the model’s ability to understand and generate core functional logic based on the specification. The RTLLM benchmark consists of four categories: Arithmetic, Memory, Control, and Miscellaneous, covering a wide range of challenges, including FSMs, combinational and sequential circuits, and long-text specifications. It poses a higher level of difficulty compared to VerilogEval. RTLLM-V2 expands upon RTLLM-V1, increasing the total number of tasks from 29 to 50, with V1 forming a subset of V2. Since some prior works only report results on RTLLM-V1, we also include our performance on V1 for fair comparison.
@@
We adopt Qwen2.5-Coder-7B-Instruct (Hui et al. 2024) as the backbone model and follow the proposed two-stage framework, applying supervised fine-tuning (SFT) for 2 epochs with LLaMaFactory (Zheng et al. 2024) in Stage I, and CRUX-enhanced GRPO training for 1 epoch with EasyR1 (Zheng et al. 2025) in Stage II.
@@
Notably, the original VerilogEval-V2 paper requires reporting only pass@1 under two settings to highlight practical LLM evaluation: a low-temperature setting (T=0.0, top_p=0.01, n=1) and a high-temperature setting (T=0.8, top_p=0.95, n=20) (Pinckney et al. 2025). However, many prior works still adopt a strategy of selecting the best performance among Temperature = 0.2, 0.5, 0.8. To enable fair comparison, we also include pass@1 results under this strategy, denoted as T-cross. For the other benchmarks, we follow common practice and generate $n=20$ independent solutions per task to compute pass@1, pass@5, and pass@10, reporting the best performance across T=0.2, 0.5, and 0.8.
@@
Our main experimental results are presented in Table 1 and Table 2. We report the performance of Foundation Models (Liu et al. 2024a; Team 2025), General Code Models (Hui et al. 2024; Guo et al. 2024), and Verilog-Specific General Models (Liu et al. 2024b; Zhao et al. 2025a; Cui et al. 2024; Wang et al. 2025a). Given the remarkable performance of Reasoning Models in recent years, we separately group Deepseek-R1 (Guo et al. 2025) and Verilog-Specific Reasoning Models (Yang et al. 2025; Zhu et al. 2025).
@@
The main results demonstrate our QiMeng-CRUX model achieves the best pass@1 performance across all benchmarks within both the General-Purpose Code Models and Verilog-Specific General Models, especially under challenging and realistic scenarios. It significantly outperforms GPT-4o on multiple tasks, most notably on RTLLM-v2, where it achieves a pass@1 of 63.8%, surpassing the previous SOTA by 12.8%, and outperforming Deepseek-V3, reaching performance comparable with Deepseek-R1. On VerilogEval-v2, QiMeng-CRUX brings substantial improvements over previous SOTA models across nearly all metrics, such as from 58.3% to 68.0% (+9.7%) in pass@1 on the CC task, and from 49.3% to 64.7% (+15.4%) on the SR task. These results suggest that CRUX offers significant advantages in tackling complex code generation tasks and real-world design inputs.
Plus, QiMeng-CRUX demonstrates robust performance across diverse task settings and specification styles. Originating from the CodeV Datasets, QiMeng-CRUX consistently performs better across tasks with diverse input forms and exhibits stronger robustness to specification variations than CodeV2.5. It achieves balanced performance across CC and SR (68.0% in CC and 64.7% in SR), avoiding the large gap observed in CodeV2.5 (58.3% in CC and 44.8% in SR). Furthermore, QiMeng-CRUX achieves substantial improvements across all metrics on the RTLLM benchmark, highlighting the effectiveness and robustness of the CRUX methodology.